# Multilevel lattice codes from Hurwitz quaternion integers

### v2

@@ VI Conclusion and Perspectives
In this work we have extended the Construction $\pi_{A}$ lattice to the Hurwitz integers and explored the potencial advantages offered by such approach.
@@
We exploit the multilevel structure of our construction to illustrate its benefits in terms of computational complexity. We have shown in Theorem 2 that it is feasible to obtain an isomorphism that guarantees this extension by splitting a two-sided prime ideal into two left-prime ideals, which induces the more complete “factorization” given by Theorem 3 and therefore the proposed Construction $\pi_{A}$. Furthermore, we highlight the applicability of this decoding process in lattice codes with high cardinality and also demonstrate that there exists a Poltyrev-good sequence of Construction $\pi_{A}$ lattices over Hurwitz integers.
@@
Some promising perspectives to be considered are the extension of this construction to other maximal orders over definite quaternion algebras where it is feasible to establish an Euclidean division algorithm, [53, 54] and to explore potential applications, such as related to the index coding problem [22].

### v3

@@ VII Conclusion and Perspectives
In this work, we have extended the Construction $\pi_{A}$ lattice to the Hurwitz integers and analyse the potential advantages offered by such an approach.
@@
We exploit the multilevel structure of our construction to illustrate its benefits in terms of computational complexity. It is shown in Theorem 2 that it is feasible to obtain an isomorphism that guarantees this extension by splitting a two-sided prime ideal into two left-prime ideals, which induces the more complete “factorization” given by Theorem 3 and therefore the proposed Construction $\pi_{A}$. The applicability of this decoding process in lattice codes with high cardinality is emphasized. In Section V we show how to construct a balanced set of codes from Construction $\pi_{A}$ for which the Averaging Lemma applies and also that there exists, with high probability, a Poltyrev-good sequence of Construction $\pi_{A}$ lattices over Hurwitz integers.
@@
Some promising perspectives to be considered are the extension of this construction to other maximal orders over definite quaternion algebras where it is feasible to establish an Euclidean division algorithm, [59, 60] and to explore potential applications, such as compute-and-forward and other properties related to the index coding problem [21, 58]. Additionally, it would be interesting to explore the possibility of generalizing the construction to the octonions using the maximal order of Octavian integers, which are closely related to the $E_{8}$ lattice.